Grade 11Magical Mathematics[Interesting Approach]

if cos(x-y) = -1 then cosx + cosy equals to what in integers
a)0
b)1
C)2
d)-1

aadi
8 Years agoGrade 11

1 Answer

aadi
8 Years ago
as cos(x-y)=-1
therefore x-y=(2n+1)pi
(x-y)/2=(2n+1)pi/2
so we can substitute odd multiples of pi/2 to verify this